OpenStudy (anonymous):

Ok so the perimeter of a quadrilateral is 44 cm. The four sides are x+7, x+6, 3x-4, and x^2-3x. How many centimeters is each side? Please help! :)

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@Jamielynne

OpenStudy (anonymous):

You add all for side teqogether and let them equal to 44 and then solve for x

OpenStudy (anonymous):

idk sorry

OpenStudy (anonymous):

but hear

OpenStudy (anonymous):

i don't get it @phuong.tong

OpenStudy (anonymous):

u add all the sides together

OpenStudy (anonymous):

(x+7) +( x+6) + (3x -4) + (X^2- 3x) = 44. Solve for x

OpenStudy (anonymous):

simplify x^2 + 2x-35 = 0

OpenStudy (anonymous):

now you factor it out and solve for x

OpenStudy (anonymous):

chose one of the x value will work which will give you 44 as perimeter. (by testing)

OpenStudy (anonymous):

um so to factor it do i have to find the greatest common factor?

OpenStudy (anonymous):

Are you there? got it?

OpenStudy (anonymous):

yeah im here

OpenStudy (anonymous):

so would that be 2x(x+1)=35 for the factorization @phuong.tong ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

After simplify you got x^2 - 2x - 35 =0 Factor this out

OpenStudy (anonymous):

(x-7) (x+5) = 0 Is it help?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

good now?

OpenStudy (anonymous):

x can not be negative

OpenStudy (anonymous):

ok thanks, but so what is the length of each side?

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

Substitu x = 7 into your equation to see if you get 44

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

nope it does not equal 44 :,( @phuong.tong

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

Sorry, my factor should be (x+7) (X-5) = 0 x= 5. Now it should work. I mess up the sign :)

OpenStudy (anonymous):

oh it must be 5

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

yes, it would give you total 44

OpenStudy (anonymous):

You are welcome. Sorry, I messed up a little